OpenStudy (anonymous):

What is the value of the expression |-20 + 5|?

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

Start with what is inside the absolute value.

OpenStudy (anonymous):

I can try to give them an example

OpenStudy (anonymous):

direct substitution: = 3(-2) + 2 -(4*-2 +5) = -6 + 2 -(-8 + 5) = -6 + 2 - (-3) = -6 + 2 +3 = -6 + 5 = -1

OpenStudy (anonymous):

I know the answer.

OpenStudy (anonymous):

but no

OpenStudy (skullpatrol):

Let the student learn please.

OpenStudy (anonymous):

sorry

OpenStudy (skullpatrol):

@emsimpson ???

OpenStudy (anonymous):

is it -15

OpenStudy (anonymous):

yay!

jimthompson5910 (jim_thompson5910):

Booisme absolute values represent distance a number is from 0 is that distance ever negative?

OpenStudy (skullpatrol):

Now take the absolute value of -15 @emsimpson

OpenStudy (skullpatrol):

$$\Huge |-15|=?$$

OpenStudy (anonymous):

that is the answer

OpenStudy (skullpatrol):

What goes where the question mark is?

OpenStudy (anonymous):

the answer is 15

OpenStudy (skullpatrol):

Correct.